HUBBLE SPACE TELESCOPE DAILY REPORT #1346 PERIOD COVERED: 0000Z (UTC) 03/13/95 - 0000Z (UTC) 03/14/95 Daily Status Report as of 073/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 5562 (WF/PC-2 Cycle 4 Calibration: Internal Monitor 3: Darks) The WF/PC-2 was used to obtain several 30-minute long, dark exposures to provide ongoing calibration of the CCD dark current and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed WF/PC-2 5142 (WF/PC-2 Imaging of Two High Redshift Galaxies Causing Absorption Systems) The WF/PC-2 was used to make observations of the high red-shifted galaxy Gal-034943 which produces absorption lines in the spectrum of QSO 0347-3819. This is to study the morphology of the absorbing galaxy. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed WF/PC-2 5655 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 2: Flats Part 2) The WF/PC-2 was used to make visual flat calibration observations. The observations were completed as planned, and no problems were reported. 1.4 Completed WF/PC-2 6143 (WF/PC-2 Cycle 4 Photometric Calibration: Smaller Set, Part II) The WF/PC-2 was used to make single CCD calibration observations of standard star GRW+70D582 to monitor the photometric stability and quantum efficiency of the PC1 and WF3. The observations were executed as scheduled, and no anomalies were reported. 1.5 Completed GHRS/2 5713 (The Distance of High Velocity Clouds) The High Resolution Spectrograph (Side Two) was used to make spectrographic observations of the Markarian galaxy MK106, which lies behind a high velocity cloud. The spectra will be centered on the Mg II doublet line at 2800 A0. A wavelength calibration observations was also taken. The observations was completed as planned, and no anomalies were completed. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 8 Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled : 11 Successful : 11 2.3 Significant HSTARs: None 2.4 Operations Notes: There were no operational problems reported. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-